Two small spheres , with a charge of Q1 on the other sphere and a charge Q2 on the other sphere ,have a distance r between their centers so thata they exert a force F on each other.What will happen to the force force F between them if one of the charges is doubles?


If bothe the charges are doubled?



1
Expert's answer
2021-06-21T16:30:29-0400

Gives

Charge "=Q_1"

Charge="Q_2"

"F=\frac{kQ_1Q_2}{r^2}=F_1"


When one charge is doubled

"Q'_1=2Q_1"

"F_2=2\frac{kQ_1 Q_2}{r^2}=2F_1"

Both force are double

"Q''_1=2Q_1"

"Q''_2=2Q_2"

"F_3=\frac{kQ''_1Q''_2}{r^2}"

"F_3=4\frac{kQ_1Q_2}{r^2}=4F_1"